OpenStudy (anonymous):

what is the solution of the equation graphed below y=-2x-3 y= 3x+2

OpenStudy (anonymous):

3x + 2 = -2x - 3 3x + 2x = -3 - 2 5x = -5 x = -1 y = -2x - 3 y = -2(-1) - 3 y = 2 - 3 y = -1 check.. y = 3x + 2 y = -2x - 3 -1 = 3(-1) + 2 -1 = -2(-1) - 3 -1 = -3 + 2 -1 = 2 - 3 -1 = -1 (correct) -1 = -1 (correct) solution is (-1,-1)

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

nope c:

OpenStudy (anonymous):

good to hear :)

OpenStudy (anonymous):

yuss (: